OpenStudy (anonymous):

PLEASE HELP what is the slope of the line that passes through the pair of points (3,8) and (9,5)? -2 2 1/2 -1/2

OpenStudy (dan815):

stop! just watch http://www.youtube.com/watch?v=WkspBxrzuZo

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given points (x1, y1) and (x2, y2) so plug in when x1=3 , y1=8 , x2=9 , y2=5

OpenStudy (anonymous):

how ? @jigglypuff314

OpenStudy (anonymous):

i need help anybody

OpenStudy (anonymous):

I got an idea.. Bettylee what is 5 - 8?

jigglypuff314 (jigglypuff314):

\[\frac{ 8-5 }{ 3 - 9 }\]

OpenStudy (anonymous):

you get 3/6 right ?

jigglypuff314 (jigglypuff314):

it would be a negative 3/6 :) which can be simplified

OpenStudy (anonymous):

how to simplify it

OpenStudy (anonymous):

Divide by 2.

OpenStudy (anonymous):

the answer is negative 1/2

OpenStudy (dan815):

yes

OpenStudy (dan815):

ill give you this.. even though you didnt try to understand, you put in a lot of effort to get the answer!!